OpenStudy (anonymous):

Can someone help me with a system of equations scenario ... IN NEED OF HELP PLEASE :'(

OpenStudy (anonymous):

sure what is the question?

OpenStudy (anonymous):

or are you supposed to make up your own?

OpenStudy (anonymous):

Part 1:For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

I need a scenario , to match those equations

OpenStudy (anonymous):

does it have to be exactly those equations for the scenario?

OpenStudy (anonymous):

Yea /: really hard.

OpenStudy (anonymous):

its really hard **

OpenStudy (anonymous):

I found one on yahoo , but its difficult . and i cant copy and paste i have to come up with my own. and im having trouble trying to figure it out.

OpenStudy (anonymous):

no problem lets try this \(3x+4y=25\) so say \(x\) is the number of budweisers solve for $3 a piece, and \(y\) is the number of heinekens sold for $4 each and the total amount of money you get is $25

OpenStudy (anonymous):

then \(2x+y=10\) says twice the number of budweisers plus the number of heinekens sold is 10

OpenStudy (anonymous):

how many of each were sold?

OpenStudy (anonymous):

the solution is \(x+3\) and \(y=4\) so you sold 3 buds and four heinies

OpenStudy (anonymous):

i mean \(x=3\)

OpenStudy (anonymous):

hi statellite help me after this

OpenStudy (anonymous):

if this will annoy your teacher, turn budweiser to filet of fish and heinekens to big macs same idea

OpenStudy (anonymous):

i put coke and sprite ahah .

OpenStudy (anonymous):

must be at the ball park, that is a lot for a coke